Compule the amount in Bits account on his 60th birthday Comput tre amount in Bobs account on his 60hbithday Do not round until the nal answer. Then round up to the nearest dolla) b. The amount in Bobs account on his 60th birthaday in S (Use the answer from part a to ind this answer Round to the nearest dollar as needed)
Answer #1

we use future value of annuity formulae to find this
FV=Pmt*[((1+r)^n)-1]/r
a)Bill pays monthly till 20 years so total 20*12=240 payments of 500 each and 9.5% interest compounded monthly
=500*(((1+(9.5%/12))^240)-1)/(9.5%/12)
=355961.77
this grows next 20 years @9.5% interest
=355961.77*(1+9.5%)^20
=2186179

b)Bob pays nothing till 40 years then for next 20 years and 20*12=240 1000 each month at 9.5% interest compounded monthly
=1000*(((1+(9.5%/12))^240)-1)/(9.5%/12)
=711923.55
same can be found using above question and it is 355961.77*2=711924